I have to find the _th term of a sequence
what i know
a= 2.2
d= -0.2
Tn= -14
how do i find what term -14 is?

Answer by waynest(281)

-14 = 2.2 + (n - 1)-0.2
-14 = 2.2 - 0.2n + 0.2
-14 = 2.4 - 0.2n
-14 - 2.4 = 2.4 - 2.4 - 0.2n
16.4 = -0.2n
-16.4/-0.2 = -0.2n/-0.2
82 = n
check:
tn = 2.2 + (82 -1)-0.2
tn = 2.2 + (81)-0.2
tn = 2.2 - 16.2
tn = -14